OpenStudy (anonymous):

What is the factored form of the following expression? w2 + 18w + 77 A. (w – 7)(w + 11) B. (w – 7)(w – 11) C. (w + 7)(w + 11) D. (w + 1)(w + 77)

OpenStudy (anonymous):

explain how to do it please.

OpenStudy (callisto):

w^2 + 18w + 77 = w^2 + 7w +11w + 77 = w(w+7) + 11(w+7) = ...??

OpenStudy (anonymous):

this questions are not hard, all you have to do is to multiple them. Start off by eliminating the ways that are not the possible answers.

OpenStudy (anonymous):

just don't understand this class

OpenStudy (anonymous):

how do u get 7w and 11w

OpenStudy (anonymous):

ok, look at the last number its +77, right?? so eliminate A and B because it will be -77

OpenStudy (anonymous):

ok

OpenStudy (callisto):

since 7w + 11w = 18w and w^2 + 7w = w(w+7) while 11w + 77 = 11 (w+7)

OpenStudy (anonymous):

continue please

OpenStudy (callisto):

Special remainder: First, consider the last term ''+ 77 '', it tells us that the factors are either negative for both or positive for both. Since '-' x '-' = '+' and '+' x '+' = '+' Then, consider the second term '+18w', it tells us that the factors are both positive. Since when you expand it, only both positive can give you a positive middle term as a whole. Take an example: (x+1)(x+2) = x^2 + x+2x +2 = x^2 + 3x +2 (x-1)(x-2) = x^2 - x - 2x +2 = x^2 - 3x +2

OpenStudy (anonymous):

so the answer is C. (w+7)(w+11)

OpenStudy (callisto):

Back to the question, since you get this step: = w(w+7) + 11(w+7) Take out the common factor and group the rest. Yes - the answer is C

OpenStudy (anonymous):

C. (w+7)(w+11) I got this correct on my test.